OpenStudy (griffindore):

What is the value of the expression 3.6 x 10^8 /1.2 x 10^3? 2.4 ⋅ 10^5 3.0 ⋅ 10^11 3.0 ⋅ 10^5 2.4 ⋅ 10^11

OpenStudy (astrophysics):

\[\frac{ 3.6 \times 10^8 }{ 1.2 \times 10^3 } \implies \frac{ 3.6 }{ 1.2 } \times \frac{ 10^8 }{ 10^3 }\] do them separately, apply exponent rules for the \[\frac{ 10^8 }{ 10^3 }\]

OpenStudy (astrophysics):

\[\huge \frac{ x^m }{ x^n } = x^{m-n}\]

OpenStudy (griffindore):

oh ok so its 10^8/10^3 = 10^8-3

OpenStudy (astrophysics):

Yes!

OpenStudy (griffindore):

which is 10^5

OpenStudy (astrophysics):

yup

OpenStudy (griffindore):

then after that its 3.6/1.2 = 3

OpenStudy (astrophysics):

Yup :)

OpenStudy (griffindore):

so its 10^5 x 3

OpenStudy (astrophysics):

3 x 10^5, but yes you got the idea!

OpenStudy (griffindore):

ya thanks